Does (1, -7) make the equation y = 5x − 7 true?








Found 2 solutions by richwmiller, math_helper:
Answer by richwmiller(17219)

Plug them in the equation y = 5x − 7 and find out.
-7 = 5(1) − 7
What do you think?

Answer by math_helper(2461)

To check, just plug in the (x, y) = (1,-7) values to see if the equation holds:
When writing I like to put a question mark above the equal sign, but here I will use =?=

-7 =?= 5(1) - 7
-7 =?= -2

NO (1, -7) does not make the equation true!

I hope you now understand how to do these.